# The State of Open Science in Software Engineering Research: A Case Study of ICSE Artifacts

### v2

@@ Abstract.
Replication packages are crucial for enabling transparency, validation, and reuse in software engineering (SE) research. While artifact sharing is now a standard practice and even expected at premier SE venues such as ICSE, the practical usability of these replication packages remains underexplored. In particular, there is a marked lack of studies that comprehensively examine the executability and reproducibility of replication packages in SE research. In this paper, we aim to fill this gap by evaluating 100 replication packages published in ICSE proceedings over the past decade (2015–2024). We assess the (1) executability of the replication packages, (2) efforts and modifications required to execute them, (3) challenges that prevent executability, and (4) reproducibility of the original findings. We spent approximately 650 person-hours in total executing the artifacts and reproducing the study findings. Our findings reveal that only 40% of the 100 evaluated artifacts were executable, of which 32.5% (13 out of 40) ran without any modification. Regarding effort levels, 17.5% (7 out of 40) required low effort, whereas 82.5% (33 out of 40) required moderate to high effort to be executed successfully. We identified five common types of modifications and 13 challenges that lead to execution failure, encompassing environmental, documentation, and structural issues. Among the executable artifacts, only 35% (14 out of 40) reproduced the original results. These findings highlight a notable gap between artifact availability, executability, and reproducibility. Our study proposes three actionable guidelines to improve the preparation, documentation, and review of research artifacts, thereby strengthening the rigor and sustainability of open science practices in SE research.
@@
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazilisbn: 979-8-4007-2025-3/2026/04ccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Empirical software engineeringccs: Software and its engineering Reusabilityccs: General and reference Evaluation
@@ 1. Introduction
Open science promotes transparency, accessibility, and reusability by encouraging researchers to share publications, data, code, and methods with the community (Mendez et al., 2020). Backed by mandates from governments, funding organizations, and publishers (OECD, 2015; The Royal Society, 2012), it has transformed fields like the natural and social sciences by enhancing reproducibility, reducing redundancy, and accelerating progress (Mirowski, 2018; Pontika et al., 2015). Although software engineering (SE) research is highly empirical and tool-centric, making it especially well-suited for open science practices, the adoption of open science in SE has been slow and inconsistent (Méndez Fernández et al., 2019). Nevertheless, in recent years, leading SE conferences and journals have taken steps to promote open science by introducing policies and artifact evaluation tracks, offering dedicated reviews and badges to incentivize the sharing of data, tools, and code (Ernst et al., 2021). A recent study shows encouraging trends: 67.7% of SE papers from top-tier conferences (e.g., ICSE, FSE, ASE) between 2017 and 2022 included artifacts (Liu et al., 2024). However, their actual usability in terms of executability and reproducibility, which are core aspects of open science, remains uncertain (Nong et al., 2023). When curated effectively, these artifacts can enhance trust, enable reuse, and support rigorous validation and comparison across studies (Méndez Fernández et al., 2019). Thus, a systematic investigation into the state of open science in SE is warranted to assess current practices, identify challenges to executability and reproducibility, and guide future improvements.
@@
Several studies have explored open science in SE through the inspection of artifact metadata, surveys, literature reviews, and methodological analyses (Solari et al., 2018; Cordeiro and Oliveira Jr, 2025a; Mahmood et al., 2018; Gundersen and Kjensmo, 2018b). For example, Solari et al. (Solari et al., 2018) analyzed 26 lab packages to propose a reference model for artifact structure, highlighting common organizational patterns. Cordeiro et al. (Cordeiro and Oliveira Jr, 2025a) surveyed SE researchers to understand reproducibility practices in controlled experiments. Rodríguez-Pérez et al. (Rodríguez-Pérez et al., 2018b) systematically reviewed 187 SZZ-related studies, which assessed reproducibility and reported limitations. In a broader context, Gundersen et al. (Gundersen and Kjensmo, 2018b) analyzed 400 AI papers to evaluate the availability of source code and the quality of documentation required for reproducibility. However, they did not attempt to execute or reproduce the studies. Going further, Nong et al. (Nong et al., 2023) executed 11 tools from 55 papers but focused narrowly on deep learning-based vulnerability detection in C/C++. While these efforts offer valuable insights, they primarily rely on secondary sources, with limited direct engagement in executing and evaluating artifacts. As a result, a comprehensive and hands-on assessment of open science practices across the broader SE research landscape remains an open challenge.
@@
In this study, we present the first large-scale empirical assessment of open science practices in SE by analyzing the executability, and reproducibility of 100 replication packages from International Conference on Software Engineering (ICSE) papers published over the past decade. First, we collected all ICSE research-track papers published from 2015 to 2024 and systematically inspected each paper to determine whether it provided a replication package link. For those who included a link, we visited it to see if it is still alive and accessible. We then reviewed the contents of each accessible package to determine whether it included executable components. Second, from the set of packages with executable components, we selected 100 replication packages using stratified random sampling (Koyuncu and Kadilar, 2009) and executed them in an isolated environment to systematically assess the degree of success and the effort required to achieve it. Third, for successfully executed packages, we compared the obtained results with those reported in the original studies to assess reproducibility. We spent approximately 650 person-hours executing the replication packages and analyzing their reproducibility. In particular, we made four major contributions by answering the following four research questions.
@@
$\bullet$ RQ1 (Executability, Effort & Modification): To what extent are replication packages executable, and what level of efforts and modifications is required to execute them? Executability is essential for enabling validation, reuse, and extension of research findings. Non-functional artifacts offer limited value and undermine scientific progress. We executed 100 ICSE replication packages in a controlled environment, strictly following the documentation provided by the original authors. Execution outcomes were labeled as executable, partially executable, or not executable. For each artifact, we recorded the required effort, categorized as low, moderate, or high, and grouped the necessary modifications into five dimensions (e.g., environment setup, code updates).
@@
$\bullet$ RQ2 (Types of Modifications): What types of modifications are needed to execute replication packages? Identifying the nature and frequency of required modifications reveals recurring quality issues in artifact packaging and helps guide best practices. We conducted a qualitative analysis and documented all modifications during execution. These were organized into a hierarchical taxonomy comprising of five broad categories and 14 specific types.
@@
$\bullet$ RQ3 (Execution Barriers): What challenges prevent the successful execution of replication packages? Identifying common failure points helps improve artifact design and review standards. For artifacts that failed or were partially executable, we documented the challenges that prevented execution and categorized them into three broader themes (e.g., documentation gaps, environment setup) and 13 specific types, along with their frequencies.
@@
$\bullet$ RQ4 (Reproducibility): Do executable replication packages reproduce the original results, and what factors contribute to reproducibility failure? Reproducibility is critical for verifying empirical results and ensuring the credibility of research claims. We evaluated 64 artifacts that were either fully or partially executable and categorized reproduction outcomes as fully reproducible, partially reproducible, not reproducible, unverified, and no output generated. We also examined factors that contribute to reproducibility failures.
@@
Full Replication Package of our study is available in our online appendix (Replication Package, 2026).
@@
Figure 1. Example of an issue in a replication package, where users encountered critical execution errors.
@@ 2. Motivating Example
Despite growing efforts to promote open science, a notable gap exists between the availability of replication packages and their practical usability (e.g., executability and reproducibility). Consider two real-world examples that illustrate this gap. Figure 1 illustrates an executability issue in a replication package. In particular, it highlights an open issue (SeamFuzz Contributors, 2024) from an ICSE 2023 paper (Lee et al., 2023), in which a user attempted to run the shared artifact but encountered critical errors that may have been caused by missing files and/or undocumented environmental assumptions. Notably, this artifact earned both the available and reusable badges (Association for Computing Machinery, 2024), suggesting that current artifact evaluation criteria may not fully reflect practical usability challenges. Figure 2 presents another case (LTL-Fuzzer Contributors, 2024) from the ICSE artifact (Meng et al., 2022), where a user reported that the code was not usable out-of-the-box due to missing required updates that had not been pushed to the public repository. As a result, the replication attempt failed despite the artifact being publicly available and endorsed.
Figure 2. Example of a user’s concern about the quality of the artifact.
@@
Motivated by these observations, this study presents the first large-scale empirical assessment of 100 ICSE artifacts to examine their executability and reproducibility. Our findings uncover executability challenges, highlight best practices, and provide actionable recommendations to strengthen the state of open science in SE.
Table 1. Summary of ICSE Papers, Artifact Availability, and Executability (2015–2024)
@@ 3.1. Dataset Construction
To evaluate open science practices in SE, we constructed a curated dataset of ICSE research papers and their associated replication artifacts spanning the past ten years. This section outlines our rationale, paper collection strategy, artifact identification process, and availability assessment.
Venue Selection Rationale. Software engineering research encompasses a diverse set of subdomains and publication venues. Conducting a comprehensive evaluation of open science practices across the entire SE landscape would be infeasible. To ensure both representativeness and focus, we selected ICSE as the target venue for our study. As a premier conference in the field, ICSE consistently publishes high-quality technical research across a diverse range of SE topics, making it an ideal candidate for evaluating artifact sharing and reproducibility practices at scale.
@@
Artifact Identification Process. To determine whether a paper shared replication package links, we employed the following three-pass search strategy. In Pass 1, we searched for open science–related keywords (e.g., replication, reproduce, artifact, package, dataset) in the PDF to identify artifact links. To construct this keyword list, we manually examined an initial set of 20 papers, which resulted in a list of 25 terms (please refer to our replication package (Replication Package, 2026)) used consistently throughout the analysis. If no link was identified via keyword search, in Pass 2, we manually scanned the full text, focusing on the end of the introduction, dedicated sections (e.g., Data Availability), footnotes, and the reference list to locate any implicit mentions of artifacts or URLs. For papers that discussed replication packages but did not include explicit artifact links, in Pass 3, we conducted a web search using the paper title and author names to check for artifacts hosted on platforms such as GitHub, Zenodo, or institutional websites. Interestingly, we found one artifact link in this process. Among 1,372 ICSE papers, 1,085 (79.08%) included at least one artifact link. While some artifacts were hosted on multiple platforms (21 cases), we collected all provided links to assess their availability.
Artifact Availability with Executable Components. For papers that included artifact links, we applied two filtering criteria: (1) availability – the provided links were active, and the artifact contents could be accessed at the time of inspection, and (2) executability – the artifacts contained executable components (e.g., code, scripts, or tools) to initiate an execution attempt. Based on these filtering criteria, we identified 941 papers with available artifact links and 796 papers whose artifacts were both available and included executable components. Among 1372 papers, 287 did not provide any replication package link, and 144 provided links that were inaccessible.
@@ 3.2. Assessing Executability, Effort, and Required Modifications
To evaluate the executability of replication artifacts, we conducted a structured execution experiment involving two highly experienced professionals. The primary execution was performed by the first author, who has over 12 years of software development experience across diverse technology stacks (e.g., Java, Python, C/C++, Docker) and previously served as a Software Technical Lead in a reputed software firm. This author spent approximately 500 person-hours executing 100 replication packages. To ensure fair evaluation and reduce individual bias, the second author, with over 15 years of industry experience, independently re-attempted execution for any package where the first author was unsuccessful, contributing an additional 150 person-hours. Of the 37 artifacts that the first author could not execute, the second author re-ran all of them and was able to change the status of only one by identifying a missing execution parameter that enabled partial execution. When both evaluators failed, either partially or entirely, we classified the artifact as Partially Executable or Not Executable, respectively.
We followed a two-step procedure for executing each artifact. First, we reviewed the documentation provided with the replication package to understand its functionality and intended execution flow. Documentation formats varied widely, commonly found in README.md files, plain text files bundled in the archive, or inline explanations embedded within code files (e.g., Jupyter Notebooks). For artifacts lacking documentation, we attempted to infer the execution logic directly from the source code. We imposed a two-hour time limit to bound our effort in such cases. Once an execution strategy was identified, we proceeded with artifact setup and execution in a controlled local environment. The following subsections detail the specific steps performed during the execution process.
@@ 3.2.2. Setting up Isolated Execution Environments
To ensure consistent and replicable execution, we prepared isolated environments according to the needs of each artifact. The details are described below.
$\bullet$ Containerization and Virtual Environments. We used Docker containers to simulate specific operating system (OS) versions and complex dependency stacks, particularly for artifacts requiring older Linux distributions. Conda environments were employed to manage Python-related dependencies in a clean and reproducible manner. While several artifacts incorporated their isolated environments (e.g., Dockerfiles or environment files), others required us to provision custom Docker images based on the documented requirements. This approach allowed us to emulate legacy setups without altering the host system and ensured consistency across executions. For example, one artifact from 2015 required Ubuntu 12.04. As getting a machine with Ubuntu 12.04 in 2025 was impossible, we successfully replicated the environment using Docker.
$\bullet$ Handling Missing Specifications. Our primary source for system specifications was the artifact’s documentation. However, when artifacts did not specify system requirements, we used our M2 machine (Table 2) as the default hardware configuration. To approximate the OS version, we applied a two-year backward buffer from the artifact’s publication year, considering the typical gap between the start of the experiment and the publication date. For example, for an artifact published in 2020, we used the latest long-term support OS version available in 2018.

### v4

@@ Abstract.
Replication packages are crucial for enabling transparency, validation, and reuse in software engineering (SE) research. While artifact sharing is now a standard practice and even expected at premier SE venues such as ICSE, the practical usability of these replication packages remain underexplored. In particular, there is a marked lack of studies that comprehensively examine the executability and reproducibility of replication packages in SE research. In this paper, we aim to fill this gap by evaluating 100 replication packages published in ICSE proceedings over the past decade (2015–2024). We assess the (1) executability of the replication packages, (2) efforts and modifications required to execute them, (3) challenges that prevent executability, and (4) reproducibility of the original findings for those that are executable. We spent approximately 650 person-hours in total to execute the artifacts and reproduce the study findings. Our analysis shows that only 40 of the 100 evaluated artifacts were fully executable. Among these, 32.5% ran without any modification. However, even executable artifacts required varying levels of effort: 17.5% required low effort, while 82.5% required moderate to high effort to execute successfully. We identified five common types of modifications and 13 challenges that lead to execution failure, encompassing environmental, documentation, and structural issues. Among the executable artifacts, only 35% (14 out of 40) reproduced the original results. These findings highlight a notable gap between artifact availability, executability, and reproducibility. Our study proposes three actionable guidelines to improve the preparation, documentation, and review of research artifacts, thereby strengthening the rigor and sustainability of open science practices in SE research.
@@
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3787813isbn: 979-8-4007-2025-3/2026/04ccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Empirical software engineeringccs: Software and its engineering Reusabilityccs: General and reference Evaluation
@@ 1. Introduction
Open science promotes transparency, accessibility, and reusability by encouraging researchers to share publications, data, code, and methods with the community (Mendez et al., 2020). Backed by mandates from governments, funding organizations, and publishers, open science has gained strong policy and institutional support and is widely promoted for enabling reproducibility, transparency, and more efficient research practices (OECD, 2015; The Royal Society, 2012; Mirowski, 2018; Pontika et al., 2015). Software engineering (SE) research is highly empirical and tool-centric, making it naturally well-suited for open science practices. However, the adoption of open science in SE has remained slow and inconsistent (Mendez et al., 2020). In recent years, leading SE conferences and journals have taken steps to promote open science by introducing policies and artifact evaluation tracks, offering dedicated reviews and badges to incentivize the sharing of data, tools, and code (Timperley et al., 2021). A recent study shows encouraging trends: 67.7% of SE papers from top-tier conferences (e.g., ICSE, FSE, ASE) between 2017 and 2022 included artifacts (Liu et al., 2024). However, their actual usability in terms of executability and reproducibility, which are core aspects of open science, remains uncertain. When curated effectively, these artifacts can enhance trust, enable reuse, and support rigorous validation and comparison across studies. Thus, a systematic investigation of the state of open science in SE is warranted to assess current practices, identify challenges to executability and reproducibility, and guide future improvements.
@@
Several studies have explored open science in SE by examining artifact metadata, conducting surveys, reviewing the literature, and performing methodological analyses (Solari et al., 2018; Cordeiro and Oliveira Jr, 2025a; Mahmood et al., 2018; Gundersen and Kjensmo, 2018). For example, Solari et al. (2018) analyzed seven lab packages and proposed a reference model for artifact structure, highlighting common structural patterns. Cordeiro and Oliveira Jr (2025a) surveyed SE researchers to understand reproducibility practices in controlled experiments. Rodríguez-Pérez et al. (2018) systematically reviewed 187 SZZ-related studies, which assessed reproducibility and reported their limitations. In a broader context, Gundersen and Kjensmo (2018) analyzed 400 AI papers to evaluate the availability of source code and the quality of documentation required for reproducibility. However, they did not attempt to execute or reproduce the studies. Furthermore, Nong et al. (2022) evaluated 11 tools across 55 papers, but focused narrowly on deep learning-based vulnerability detection in C/C++. While these efforts offer valuable insights, they primarily rely on secondary sources, with limited direct engagement in the execution and evaluation of artifacts. Therefore, a comprehensive and hands-on assessment of open science practices across the broader SE research landscape remains an open challenge.
@@
In this study, we present a comprehensive empirical assessment of open science practices in SE by analyzing the executability and reproducibility of 100 replication packages from International Conference on Software Engineering (ICSE) papers published over the past decade. First, we collected all ICSE research-track papers published from 2015 to 2024 and systematically inspected each paper to determine whether it included a link to a replication package. For those who included a link, we visited it to see if it is still alive and accessible. We then reviewed the contents of each accessible package to determine whether it included executable components. Second, from the set of packages with executable components, we selected 100 replication packages using stratified random sampling (Koyuncu and Kadilar, 2009) and executed them in an isolated environment to systematically assess the degree of success and the effort required to achieve it. Third, for successfully executed packages, we compared the obtained results with those reported in the original studies to assess reproducibility. We spent approximately 650 person-hours executing the replication packages and analyzing their reproducibility. In particular, we made four major contributions by answering the following four research questions.
@@
$\bullet$ RQ1 (Executability, Effort & Modification): To what extent are replication packages executable, and what level of effort and modifications are required to execute them? Executability is essential for validating, reusing, and extending research findings. Non-functional artifacts offer limited value and undermine scientific progress. We executed 100 ICSE replication packages in a controlled environment, strictly following the original authors’ documentation. Execution outcomes were labeled as executable, partially executable, or not executable. For each artifact, we recorded the required effort, categorized as low, moderate, or high, and grouped the necessary modifications into five dimensions (e.g., environment setup, code updates).
@@
$\bullet$ RQ2 (Types of Modifications): What types of modifications are needed to execute replication packages? Identifying the nature and frequency of required modifications reveals recurring quality issues in artifact packaging and helps guide best practices. We thus attempt to execute replication packages manually and document all required modifications. These were organized into a hierarchical taxonomy comprising five broad categories and 14 specific types.
@@
$\bullet$ RQ3 (Execution Barriers): What challenges prevent the successful execution of replication packages? Identifying common failure points helps improve artifact design and review standards. For artifacts that failed or were partially executable, we documented the challenges that prevented execution and categorized them into three broader themes (e.g., documentation gaps) and 13 specific types, along with their frequencies.
@@
$\bullet$ RQ4 (Reproducibility): Do executable replication packages reproduce the original results, and what factors contribute to reproducibility failure? Reproducibility is critical for verifying empirical results and ensuring the credibility of research claims. We evaluated 64 artifacts that were either fully or partially executable and categorized reproduction outcomes as fully reproducible, partially reproducible, not reproducible, unverifiable, and no output generated. We also examined factors that contribute to reproducibility failures.
@@
Full Replication Package of our study can be found in our online appendix (Muttakin et al., 2026).
@@
Figure 1. Example of an issue (SeamFuzz Contributors, 2024) in a replication package, where users encountered critical execution errors.
@@ 2. Motivating Example
Despite growing efforts to promote open science, a notable gap exists between the availability of replication packages and their practical usability (e.g., executability and reproducibility). Consider two real-world examples that illustrate this gap. Figure 1 illustrates an executability issue in a replication package. In particular, it highlights an issue reported in the replication package uploaded to GitHub for an ICSE 2023 paper (Lee et al., 2023), where a user attempted to run the shared artifact but encountered critical errors, potentially due to missing files or undocumented environmental assumptions. Notably, this artifact received both the available and reusable badges (Association for Computing Machinery, 2024), indicating that existing artifact evaluation criteria may not always capture certain practical usability challenges. Figure 2 presents another case from the ICSE artifact (Meng et al., 2022), where a user reported that the code was not usable out-of-the-box due to missing required updates that had not been pushed to the public repository. As a result, the replication attempt was unsuccessful despite the artifact being publicly available and endorsed.
Figure 2. Example of a user’s concern about the quality of the artifact (LTL-Fuzzer Contributors, 2024).
@@
Motivated by these observations, we present a large-scale empirical assessment of 100 ICSE artifacts to examine their executability and reproducibility. Our findings uncover executability challenges, highlight best practices, and provide actionable recommendations to strengthen the state of open science in SE.
Table 1. Summary of ICSE Papers, Artifact Availability, and Executability (2015-2024)
@@ 3.1. Dataset Construction
To evaluate open science practices in SE, we constructed a curated dataset of ICSE research papers and their associated replication artifacts spanning the past ten years (2015–2024). The following sections outline our rationale, paper collection strategy, artifact identification process, and availability assessment.
Venue Selection Rationale. Software engineering research encompasses a diverse set of subdomains and publication venues. Conducting a comprehensive evaluation of open science practices across the entire SE landscape would be infeasible. To ensure both representativeness and focus, we selected ICSE as the target venue for our study. As a premier conference in the field, ICSE consistently publishes high-quality technical research across a diverse range of SE topics, making it an ideal venue for evaluating artifacts, their executability, and reproducibility at scale.
@@
Artifact Identification Process. To determine whether a paper shared replication package links, we employed the following three-pass search strategy. In Pass 1, we searched the PDF for open-science-related keywords (e.g., replication, reproduce, artifact, package, dataset) to identify artifact links. To construct this keyword list, we manually examined an initial set of 20 papers, which resulted in a list of 25 terms (please refer to our replication package (Muttakin et al., 2026)) used consistently throughout the analysis. If no link was identified via keyword search, in Pass 2, we manually scanned the full text, focusing on the end of the introduction, dedicated sections (e.g., Data Availability), footnotes, and the reference list to locate any implicit mentions of artifacts or URLs. For papers that discussed replication packages but did not include explicit artifact links, in Pass 3 we conducted web searches using the paper title and author names to determine whether artifacts were available on platforms such as GitHub, Zenodo, or institutional websites. Interestingly, we found one artifact link in this process. Among 1,372 ICSE papers, 1,085 (79.08%) included at least one artifact link. While some artifacts were hosted on multiple platforms (21 cases), we collected all provided links to assess their availability.
Artifact Availability with Executable Components. For papers that included artifact links, we applied two filtering criteria: (1) availability – the provided links were active, and the artifact contents could be accessed at the time of inspection, and (2) executability – the artifacts contained executable components (e.g., code, scripts, or tools) to initiate an execution attempt. Based on these filtering criteria, we identified 941 papers with available artifact links and 796 papers whose artifacts were truly available and had executable components. Among 1372 papers, 287 did not provide any replication package link, and 144 provided links that were inaccessible.
@@ 3.2. Assessing Executability, Effort, and Required Modifications
To evaluate the executability of replication artifacts, we conducted a structured execution experiment involving two highly experienced professionals. The primary execution was conducted by the first author, who has over 12 years of software development experience across diverse technology stacks (e.g., Java, Python, C/C++, Docker) and previously served as a Software Technical Lead at a reputable software firm. This author spent approximately 500 person-hours executing 100 replication packages. To ensure fair evaluation and reduce individual bias, the second author, with over 15 years of development experience, independently re-attempted execution for any package where the first author was unsuccessful, contributing an additional 150 person-hours. Of the 37 artifacts that the first author could not execute, the second author re-ran all of them and changed the status of only one by identifying a missing execution parameter that enabled partial execution. When both evaluators failed, either partially or entirely, we classified the artifact as Partially Executable or Not Executable, respectively.
We followed a two-step procedure for executing each artifact. First, we reviewed the documentation provided with the replication package to understand its functionality and intended execution flow. Documentation formats varied widely and were commonly provided through README.md files, plain text files bundled in the archive, or inline explanations embedded within code files (e.g., Jupyter Notebooks). For artifacts lacking documentation, we attempted to infer the execution logic directly from the source code. We imposed a two-hour time limit to bound our effort in such cases. Once an execution strategy was identified, we proceeded with artifact setup and execution in a controlled local environment. The following sections detail the steps performed during execution.
@@ 3.2.2. Setting up Isolated Execution Environments
To ensure consistent and replicable execution, we prepared isolated environments tailored to each artifact’s requirements, as discussed below.
$\bullet$ Containerization and Virtual Environments. We used Docker containers to simulate specific operating system (OS) versions and complex dependency stacks, particularly for artifacts requiring older Linux distributions. Conda environments were used to manage Python-related dependencies in a clean, reproducible manner. While several artifacts used their isolated environments (e.g., Dockerfiles or environment files), others required us to provision custom Docker images according to the documented requirements. This approach allowed us to emulate legacy setups without altering the host system and ensured consistency across executions. For example, one artifact from 2015 required Ubuntu 12.04. Since obtaining a machine running Ubuntu 12.04 in 2025 was impossible, we successfully replicated the environment using Docker.
$\bullet$ Handling Missing Specifications. Our primary source for system specifications was the artifact’s documentation. However, when artifacts did not specify system requirements, we used our M2 machine (Table 2) as the default hardware configuration. To approximate the OS version, we used a two-year backward buffer from the artifact’s publication year, accounting for the typical gap between the start of the experiment and the publication date. For example, for an artifact published in 2020, we used the latest long-term support OS version available in 2018.